OpenStudy (anonymous):

Check my answer. Question: Write the equation that describes the line in slope-intercept form.slope = 4, point (3, –2) is on the line. My answer: So first we use the point slope form. y - y1 = m (x - x1) = y - 3 = 4 (x - (- 2) = y - 3 = 4x + 8. Now we add 3 to both sides y - 3 + 3 = 4x + 8 + 3 = y = 4x + 11.

OpenStudy (anonymous):

You accidenatlly used x, as y and vice versa

OpenStudy (anonymous):

My bad. Is this correct now? y - ( -2) = 4 (x - 3) = y + 2 = 4x - 12. Now we subtract 2 from both sides y + 2 - 2 = 4x - 12 - 2 = y = 4x - 13.

OpenStudy (anonymous):

Not quite. -12-2=14. but other than that. You are right.

OpenStudy (anonymous):

i meant -14 btw

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Anytime